Q)Ram got an order from a garment manufacturer for 480 Denim Shirts. He bought 12 sewing machines and appointed some tailors to do the job. However, many didn't report for duty. As a result,each had to stitch 32 more shirts than originally planned by Ram , with equal distribution of work.How many tailors had been appointed earlier and how many had not reported for work?
(a)12,4 (b)10,3 (c)10,4 (d) none of these

Answer by jorel1380(2518)

480/10=48
48+32=80
480/80=6
10 tailors were needed. Only 6 reported..